OpenStudy (anonymous):

please help!!!!! 1 medal and fan M(4, 2) is the midpoint of RS. The coordinates of S are (6, 1). What are the coordinates of R? (2, 3) (1.5, 3) (5, 1.5) (8, 0)

OpenStudy (anonymous):

I THINK ITS (2,3) BUT IDK

OpenStudy (anonymous):

4 = (6 + x)/2 and 2 = (1 + y)/2 8 = 6 + x and 4 = 1 + y x = 2 and y = 3 A

OpenStudy (anonymous):

SO IM RIGHT?

OpenStudy (anonymous):

yuo =D

OpenStudy (anonymous):

thanks :) its appreciated